A stone is thrown straight upwards and it rises to a maximum height of 20m. With what speed was it thrown?

1
Expert's answer
2022-02-18T15:53:34-0500

The maximum height

"h_{\max}=\frac{v_0^2}{2g}"

The initial speed

"v_0=\sqrt{2gh_{\max}}=\sqrt{2*9.8*20}=19.8\:\rm m/s"